Circular No. 5780

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993L IN IC 5270 Robert Evans, Hazelbrook, New South Wales, Australia, reports his

visual discovery of an apparent supernova located ∼ 40′′ west and 15′′ north of the center of IC 5270 (α = 22^{h}55^{m}.2, δ = −36^{o}07′, equinox 1950.0). Upon first detection at Apr. 30.8 UT, he estimated m_{v} ∼ 14. R. H. McNaught, Anglo-Australian Observatory, reports the following precise position from a short exposure taken in twilight on Apr. 30.83 with the Uppsala Southern Schmidt telescope: α = 22^{h}55^{m}05^{s}.81, δ = −36^{o}07′31′′.0 (equinox 1950.0, uncertainty 0′′.4 in each coordinate); the galaxy was only weakly recorded, but the supernova appears due west.

SUPERNOVA 1993J IN NGC 3031 S. R. Trammell, D. C. Hines, and J. C. Wheeler, University of Texas,

report: “Optical spectropolarimetry (range 420-690 nm; dispersion 0.269 nm/pixel) was obtained with the 2.7-m telescope (+ Large Cassegrain Spectrograph) at McDonald Observatory on Apr. 20.11 UT. We detect a relatively high polarization of 0.9 ± 0.1 at angle 33^{o} ± 3^{o}, and there is a 0.5-percent drop in polarization across Hα accompanied by a 10^{o} positionangle rotation. These findings are consistent with those reported by B. Jannuzzi et al. (IAUC 5776) for data taken after ours on Apr. 26.12. We measure the polarization of the Hα emission alone to be 1.1 ± 0.1 at a position angle of 150^{o} ± 4^{o}. From these data, we infer the presence of two distinct polarization components and conclude that at least part of the polarization is intrinsic to the supernova. This suggests an asymmetric geometry (see also IAUC 5577).”

J. A. Smith, Florida Institute of Technology, reports the following infrared magnitudes (± 0.02 mag unless otherwise noted) for SN 1993J, obtained using the Kitt Peak 1.3-m telescope (+ SQIID): Apr. 20.212 UT, J = 10.02, H = 10.02, K = 9.84 ± 0.03; 21.265, 10.05, 10.02, 9.85 ± 0.03; 22.165, 10.06, 10.02, 9.87 ± 0.03; 24.145, 10.14, 10.04, 9.89 ± 0.03; 25.246, 10.16 ± 0.01, 10.09, 9.89 ± 0.03. Checks were made against the two primary GSC references reported by Odewahn et al. on IAUC 5760.

Further photoelectric V photometry (cf. IAUC 5750, 5755) by M. Okyudo and T. Ishida, Nishi-Harima Astronomical Observatory, communicated by T. Kato, Kyoto University: Apr. 23.510 UT, 11.27 (bad conditions); 25.719, 11.40; 26.583, 11.55; 26.740, 11.46; 27.609, 11.64.

1993 May 1 Daniel W. E. Green